# VisionLLaMA: A Unified LLaMA Interface for Vision Tasks

### v1

@@ Abstract
Large language models are built on top of a transformer-based architecture to process textual inputs. For example, the LLaMA family of models stands out among many open-source implementations. Can the same transformer be used to process 2D images? In this paper, we answer this question by unveiling a LLaMA-like vision transformer in plain and pyramid forms, termed VisionLLaMA, which is tailored for this purpose. VisionLLaMA is a unified and generic modeling framework for solving most vision tasks. We extensively evaluate its effectiveness using typical pre-training paradigms in a good portion of downstream tasks of image perception and especially image generation. In many cases, VisionLLaMA have exhibited substantial gains over the previous state-of-the-art vision transformers. We believe that VisionLLaMA can serve as a strong new baseline model for vision generation and understanding. Our code will be released at https://github.com/Meituan-AutoML/VisionLLaMA.
Figure 1: Generated image samples by our DiT-LLaMA-XL of resolution (256, 256) with a CFG ratio of $4.0$. Best viewed on screen.
@@ 1 Introduction
Large language models have aroused great interest in the research community. One of the most influential and representative work is LLaMA [66, 67]. Many recent works have converged to this architecture and solutions for various applications are built upon the open-sourced models. Besides, we have witnessed the blooming of multimodal models, where many methods also heavily rely on LLaMA for text processing and CLIP-fashioned [51] vision transformers [22] for visual perception. Meanwhile, many endeavors [23, 73, 38] have been devoted to accelerating the inference speed and/or the memory cost of LLaMA. In a word, LLaMA is now the de facto architecture.
@@
    We investigate means to adapt VisionLLaMA to tackle common vision tasks, including image comprehension and creation (Figure VisionLLaMA: A Unified LLaMA Interface for Vision Tasks). We examine two well-known vision architecture schemes (plain and pyramid) and assess their performance under supervised and self-supervised learning scenarios. Additionally, we introduce AS2DRoPE (i.e. auto-scaled 2D RoPE), which expands rotated positional encoding from 1D to 2D and utilizes interpolation scaling to accommodate arbitrary resolutions.
@@ 2 Related Work
Vision Transformer. ViT [22] successfully applied Transformer [68] from natural language processing to the vision world and many more efficient and powerful follow-up works are induced, like DeiT [65], Swin [43], PVT [70], and Twins [12]. The pre-training paradigm has been shifted from supervised learning on large-scale categorically labeled datasets like ImageNet [19] to unsupervised learning [25], and to contrastive learning on huge amounts of image-text pairs as in CLIP [51]. DiT [50] adopts a transformer that operates on latent patches for diffusion models [60, 28], outperforming the commonly used U-Net backbone [54].
@@ 3.1 Plain Transformer
Figure 2: Our VisionLLaMA block (a) in plain Transformer and its variant block (b) in pyramid Transformer.
Our plain VisionLLaMA follows the pipeline of ViT [22] and we retain the architecture design of LLaMA as closely as possible. For an image of $H\times W$, it’s firstly transformed and flattened into $N=\frac{H\times W}{P^{2}}$ non-overlapped patches $X\in{\cal R}^{N\times C}$. Then a class token is prepended at the beginning of the sequence and the whole sequence is processed by $L$ VisionLLaMA blocks. Unlike [22], we do not add positional encodings to the input sequence since our basic block readily contains positional encoding. Specifically, the basic block differs from the standard ViT block by two components: self-attention with positional encoding (RoPE) [62] and SwiGLU activation [59]. We still utilize LayerNorm [2] instead of RMSNorm [80] since we find the former behave better through the classification experiment (see Table LABEL:tab:_ablation_norm). The basic block is illustrated in Figure 2 (a). It should be noted that directly applying 1D RoPE in vision tasks cannot well generalize to other resolutions, which is different from the training resolution. Therefore, we extend it to the 2D form. It can be formally written as,
@@ 3.2 Pyramid Transformer
It’s straightforward to apply VisionLLaMA to window-based transformers that utilize additive relative position encoding, such as Swin [43]. In this paper, we choose a stronger baseline Twins [12] to explore how to build a powerful pyramid transformer under strictly controlled settings. The original architecture of Twins exploits a conditional position encoding and interleaved local-global information exchange in the form of local and global attention. These components can be found in various transformers, which means it is not difficult to apply VisionLLaMA in other pyramid transformer variants by following our method. Note that our target is not to invent a novel pyramid vision transformer, but to show how we adapt the basic design of VisionLLaMA based on the existing ones. Therefore, we simply conform to the smallest modifications to the architecture and hyperparameters. Following the name convention of [12], the two consecutive blocks can be written as,
@@ 3.3 Training or Inference Beyond the Sequence Length
$\begin{bmatrix}\cos(i\theta_{0})&-\sin{(i\theta_{0})}&0&0&\dots&0&0&0\\ \sin(i\theta_{0})&\cos{(i\theta_{0})}&0&0&\dots&0&0&0\\ 0&0&\cos(j\theta_{0})&-\sin{(j\theta_{0})}&\dots&0&0&0\\ 0&0&\sin(j\theta_{0})&-\cos{(j\theta_{0})}&\dots&0&0&0\\ \\ 0&0&0&\dots&\cos(i\theta_{d-4})&-\sin{(i\theta_{d-4})}&0&0\\ &&&&\sin(i\theta_{d-4})&\cos{(i\theta_{d-4})}&0&0\\ 0&0&0&\dots&0&0&\cos(j\theta_{d-4})&-\sin{(j\theta_{d-4}})\\ 0&0&0&\dots&0&0&\cos(j\theta_{d-4})&-\sin{(j\theta_{d-4})}\\ \end{bmatrix}$
@@
As for the GSA under the pyramid setting, we require special treatments since we need to add positional information to the summarized keys. These sub-sampled keys are generated by abstraction on the feature maps. Without loss of generality, we use a convolution with a kernel size of $k\times k$ and stride of $k$. The coordinate of the generated key can be formulated as the average of the sampled features. We show a simple example in Figure 3.
@@ 4.1 Image Generation
We choose 250 sample steps of DDPM as [50] and show the result in Table 1. As a common practice, FID is regarded as a primary metric. We also report other secondary metrics such as sFID [47], Precision/Recall [35], and Inception Score [56]. Most experiments are controlled on 400k training steps. VisionLLaMA significantly outperforms DiT across various model sizes. We also extend the training steps of XL models to 2352k steps to evaluate whether our models have the faster convergence advantage or still behave better under the setting of longer training epochs. DiT-LLaMA-XL/2 has 0.83 lower FID [27] than DiT-XL/2, indicating that VisionLLaMA not only has better computing efficiency but higher performance than DiT. We show some generated samples in Figure VisionLLaMA: A Unified LLaMA Interface for Vision Tasks using our XL model.
@@
Table 1: Image generation comparisons using the DiT framework [50]. All the models are trained using an image resolution of 256$\times$256 with a batch size of 256. Metrics are calculated using the sampled 50k images. IS: inception score [56].
[…]
SiT-S/2 ${}^{\dagger}$ 6.06 33 400 0.0001 58.15 9.12 41.01 60.23 24.72
SiT-LLaMA-S/2 6.06 33 400 0.0001 53.90 8.78 42.98 60.36 26.74
SiT-B/2 ${}^{\dagger}$ 23.01 130 400 0.0001 35.54 6.57 52.68 64.38 42.33
SiT-LLaMA-B/2 23.01 130 400 0.0001 29.53 6.32 56.07 64.07 50.13
DiT-L/2 80.71 458 400 0.0001 23.3 - - - -
SiT-L/2 ${}^{\dagger}$ 80.71 458 400 0.0001 19.34 5.28 63.00 63.60 70.47
SiT-LLaMA-L/2 80.71 458 400 0.0001 14.32 5.17 66.39 63.64 86.85
SiT-XL/2 ${}^{\dagger}$ 118.64 675 400 0.0001 16.98 5.07 65.12 64.10 77.06
SiT-LLaMA-XL/2 118.64 675 400 0.0001 12.20 5.03 67.86 63.08 95.28
Table 2: Image generation comparisons using the SiT framework [46]. All the models are trained using an image resolution of 256$\times$256 with a global batch size of 256. Metrics are calculated using the sampled 50k images without classifier-free guidance. IS: inception score. The FID is calculated by 250 steps SDE Euler sampler. $\dagger$: reproduced result using the released code.
@@ 4.2.1 Supervised Training
Plain Vision Transformer Comparison. DeiT3 [65] is the state-of-the-art plain vision transformer, which proposes special data augmentations and performs extensive hyperparameter search to boost the performance of DeiT [64]. During the reproduction of DeiT3, we observe that it is sensitive to hyperparameters and prone to overfitting. Replacing the class token with GAP (global average pooling)[13] leads to a 0.7% top-1 accuracy drop for the DeiT3-Large model after 800 epochs of training. Therefore, we use the class token instead of GAP in the plain transformer and report the result in Table 3, where VisionLLaMA achieves a top-1 accuracy comparable to DeiT3. The detailed hyperparameter is listed in the appendix. Note that the accuracy on a single resolution does not provide comprehensive comparisons, we also evaluate the performance across different image resolutions as [13] and report the result in Table 4. As for DeiT3, we use the bicubic interpolation for the learnable positional encoding. Although these two models have comparable performance at the resolution of 224$\times$224, the gap is enlarged when the resolution is increased, which means our method generalizes better across different resolutions, which is a vital function for many downstream tasks such as object detection.
Model 160 224 256 288 512 768
DeiT3-Large [65] 83.1 84.5 84.7 84.6 82.1 76.5
VisionLLaMA-L 83.1 84.6 84.7 84.8 83.5 79.1
Pyramid Vision Transformer. We use the same architecture as Twins-SVT [12] and the detailed configuration is listed in Table 17. We remove the conditional position encoding since VisionLLaMA already contains one kind of rotary position encoding. Therefore, VisionLLaMA is a convolution-free architecture. We do not tune the hyper-parameters and directly follow the setting provided in [12]. Although it’s suboptimal, it can still achieve competitive performance. As [12, 13], we do not use the class token and apply GAP. In particular, all the models are trained for 300 epochs with a batch size of 1024. The learning rate is initialized to be 0.001 and decayed to zero within 300 epochs following the cosine strategy. The result is shown in Table 3 and our method achieves comparable performance as Twins across various levels of models and outperforms Swin [43] consistently. We further compare the pyramid transformers using popular downstream tasks, which are shown in the later sections.
@@ 4.2.2 Self-Supervised Training
There are two common approaches to evaluating the performance of the self-supervised vision transformers [25] using the ImageNet dataset. In this section, we make comparisons based on these two ways. To make fair comparisons, we limit the training data to ImageNet-1K. We also exclude any component that utilizes CLIP [51], DALLE [52], or distillation, which can be orthogonally combined to further boost the performance. Our implementation is based on the MMPretrain framework [15]. We utilize the MAE framework and replace the encoder using VisionLLaMA while keeping other components unchanged. This minor modified setting forms a controlled experiment to evaluate the role of our approaches. Moreover, we use the same hyperparameter as [25], which is suboptimal to our method. Fortunately, this simple setting still achieves a significant performance boost over the strong baseline.
Full fine-tuning. In such a setting, the model is first initialized using the pre-trained weights and then trained for extra epochs with totally trainable parameters. Trained by 800 epochs on the ImageNet, VisionLLaMA-Base achieves 84.0% top-1 accuracy, which exceeds ViT-Base by 0.8%. Note that our method uses a mask ratio of 0.75 as [25], whose training speed is about 3 times faster than SimMIM [75]. We also increased the training epochs to 1600 to verify whether VisionLLaMA keeps the advantage given sufficient training resources. VisionLLaMA-Base achieves new state-of-art result among MAE variants, 84.3% top-1 accuracy, which outperforms ViT-Base by 0.9%. This result is even higher than MaskFeat [71] where new training objectives are proposed. Regarding full fine-tuning having a risk of performance saturation [42, 69], our boost is significant. Next we resort to the linear probing metric to provide extra evaluations, which is considered a more reliable evaluation for representative learning by a recent work [9].
Models Pretrain Epochs SFT Acc LP Acc
${}^{\dagger}$ [25] 800 83.2 65.1
@@
Table 5: Comparison with masked image modeling SSL methods on the ImageNet validation set. $\dagger$: reproduced in MMPretrain.
@@ 4.3.1 Supervised Training
Following [12, 43], we evaluate our method using semantic segmentation on the ADE20K [82] dataset. To make fair comparisons, we limit the baselines to only using ImageNet-1K in the pre-training stage. Specifically, we make use of the UperNet [74] framework and replace the backbone with pyramid VisionLLaMA. Our implementation is based on the MMSegmentation framework[14]. Our models are trained for 160k steps with a global batch size of 16. The detailed setting of the hyperparameter is shown in Section B.7. We report the result in Table 6. Under similar FLOPs, our method outperforms both Swin and Twins by more than 1.2% mIoU.
Models Param mIoU
@@
Table 6: Performance comparisons with different backbones on ADE20K validation dataset. All backbones are pre-trained on ImageNet-1K with labels. mIoU is evaluated by the single scale setting.
4.3.2 Self-Supervised Training
We use the UperNet [74] framework to perform semantic segmentation on the ADE20K dataset, which is a popular benchmark for backbones. We carefully control the experiment and replace the ViT backbone with VisionLLaMA while keeping other components and hyperparameters unchanged. Our implementation is based on MMSegmentation [14] and the detailed hyperparameters are provided in Section B.6. The result is given in Table 7. As for the 800 epoch pre-training groups, VisionLLaMA-B significantly boosts ViT-Base by 2.8% mIoU. It also outperforms some other modifications such as introducing extra training objectives or features [42, 71] by clear margins. Moreover, those approaches introduce extra overhead for the training process and slow down the training speed. We emphasize that the training speed of a method is becoming more and more important in the age of large models. In contrast, VisionLLaMA only involves the replacement of the base model and has the same fast training speed as [25]. In principle, our method can be seamlessly combined with these modifications. We further evaluate the performance of longer pre-training epochs of 1600, VisionLLaMA-B achieves 50.2% mIoU on the ADE20K validation set, which boosts ViT-B by 2.1% mIoU.
Models Pretrain Epochs mIoU
${}^{\dagger}$ 800 46.2
@@ 4.3.2 Self-Supervised Training
Table 7: Performance comparisons with different SSL trained backbones on ADE20K validation dataset. All backbones are pre-trained on ImageNet-1K without labels. mIoU is evaluated by the single scale setting. $\dagger$: reproduce result using [14].

### v2

# VisionLLaMA: A Unified LLaMA Backbone for Vision Tasks

@@ Abstract
We all know that large language models are built on top of a transformer-based architecture to process textual inputs. For example, the LLaMA family of models stands out among many open-source implementations. Can the same transformer be used to process 2D images? In this paper, we answer this question by unveiling a LLaMA-like vision transformer in plain and pyramid forms, termed VisionLLaMA, which is tailored for this purpose. VisionLLaMA is a unified and generic modeling framework for solving most vision tasks. We extensively evaluate its effectiveness using typical pre-training paradigms in a good portion of downstream tasks of image perception and especially image generation. In many cases, VisionLLaMA has exhibited substantial gains over the previous state-of-the-art vision transformers. It is our hope that researchers in computer vision can apply VisionLLaMA presented here to solve various specific image generation and perception tasks.
Code is at: https://github.com/Meituan-AutoML/VisionLLaMA
Keywords:
LLaMA Diffusion Model Vision Transformer
Figure 1: Generated images by DiT-LLaMA-XL of resolution (256, 256) with CFG.
@@ 1 Introduction
Large language models have aroused great interest in the research community. One of the most influential and representative work is LLaMA [64, 65]. Many recent works have converged to this architecture and solutions for various applications are built upon the open-sourced models. Besides, we have witnessed the blooming of multimodal models, where many methods also heavily rely on LLaMA for text processing. Meanwhile, many endeavors [23, 71, 37] have been devoted to accelerating the inference speed and/or the memory cost of LLaMA. In a word, LLaMA is now the de facto architecture.
@@
    We investigate means to adapt VisionLLaMA to tackle common vision tasks, including image comprehension and creation (Figure 1). We examine two well-known vision architecture schemes (plain and pyramid) and assess their performance under supervised and self-supervised learning scenarios. Additionally, we introduce AS2DRoPE (i.e. auto-scaled 2D RoPE), which expands rotated positional encoding from 1D to 2D and utilizes interpolation scaling to accommodate arbitrary resolutions.
@@ 2 Related Work
Vision Transformer. ViT [22] successfully applied Transformer [66] from natural language processing to the vision world and many more efficient and powerful follow-up works are induced, like DeiT [63], Swin [42], PVT [68], and Twins [12]. The pre-training paradigm has been shifted from supervised learning on large-scale categorically labeled datasets like ImageNet [19] to unsupervised learning [25]. DiT [49] adopts a transformer that operates on latent patches for diffusion models [58, 28], outperforming the commonly used U-Net backbone [53].
@@ 3.1 Plain Transformer
Figure 2: VisionLLaMA block (a) in plain Transformer and (b) in pyramid Transformer.
Our plain VisionLLaMA follows the pipeline of ViT [22] and we retain the architecture design of LLaMA as closely as possible. For an image of $H\times W$, it’s firstly transformed and flattened into $N=\frac{H\times W}{P^{2}}$ non-overlapped patches $X\in{\cal R}^{N\times C}$. Then a class token is prepended at the beginning of the sequence and the whole sequence is processed by $L$ VisionLLaMA blocks. Unlike [22], we do not add positional encodings to the input sequence since our basic block readily contains positional encoding. Specifically, the basic block differs from the standard ViT block by two components: self-attention with positional encoding (RoPE) [60] and SwiGLU activation [57]. We still utilize LayerNorm [3] instead of RMSNorm [78] since we find the former behave better through the classification experiment (see Table 7(b)). The basic block is illustrated in Figure 2 (a). It should be noted that directly applying 1D RoPE in vision tasks cannot well generalize to other resolutions, which is different from the training resolution. Therefore, we extend it to the 2D form. It can be formally written as,
@@ 3.2 Pyramid Transformer
It is straightforward to apply VisionLLaMA to window-based transformers that utilize additive relative position encoding, such as Swin [42]. In this paper, we choose a stronger baseline Twins [12] to explore how to build a powerful pyramid transformer under strictly controlled settings. The original architecture of Twins exploits a conditional position encoding and interleaved local-global information exchange in the form of local and global attention. These components can be found in various transformers, which means it is not difficult to apply VisionLLaMA in other pyramid transformer variants by following our method. Note that our target is not to invent a novel pyramid vision transformer, but to show how we adapt the basic design of VisionLLaMA based on the existing ones. Therefore, we simply conform to the smallest modifications to the architecture and hyperparameters. Following the name convention of [12], the two consecutive blocks can be written as,
@@ 3.3 Training or Inference Beyond the Sequence Length
$\begin{bmatrix}\cos(i\theta_{0})&-\sin{(i\theta_{0})}&0&0&\dots&0&0&0\\ \sin(i\theta_{0})&\cos{(i\theta_{0})}&0&0&\dots&0&0&0\\ 0&0&\cos(j\theta_{0})&-\sin{(j\theta_{0})}&\dots&0&0&0\\ 0&0&\sin(j\theta_{0})&\cos{(j\theta_{0})}&\dots&0&0&0\\ \\ 0&0&0&\dots&\cos(i\theta_{d-4})&-\sin{(i\theta_{d-4})}&0&0\\ &&&&\sin(i\theta_{d-4})&\cos{(i\theta_{d-4})}&0&0\\ 0&0&0&\dots&0&0&\cos(j\theta_{d-4})&-\sin{(j\theta_{d-4}})\\ 0&0&0&\dots&0&0&\cos(j\theta_{d-4})&\cos{(j\theta_{d-4})}\\ \end{bmatrix}$
@@
As for the GSA under the pyramid setting, we require special treatments since we need to add positional information to the summarized keys. These sub-sampled keys are generated by abstraction on the feature maps. Without loss of generality, we use a convolution with a kernel size of $k\times k$ and stride of $k$. The coordinate of the generated key can be formulated as the average of the sampled features. We show a simple example in Figure 5 (appendix.).
@@ 4.1 Image Generation
We choose 250 sample steps of DDPM as [49] and show the result in Table 1. As a common practice, FID is regarded as a primary metric. We also report other secondary metrics such as sFID [46], Precision/Recall [34], and Inception Score [55]. Most experiments are controlled on 400k training steps. VisionLLaMA significantly outperforms DiT across various model sizes. We also extend the training steps of XL models to 2352k steps to evaluate whether our models have the faster convergence advantage or still behave better under the setting of longer training epochs. DiT-LLaMA-XL/2 has 0.83 lower FID [27] than DiT-XL/2, indicating that VisionLLaMA not only has better computing efficiency but higher performance than DiT. We show some generated samples in Figure 1 using our XL model.
Table 1: Image generation comparisons using DiT [49].
@@ 4.2.1 Supervised Training
Plain Vision Transformer Comparison. DeiT3 [63] is the state-of-the-art plain vision transformer, which proposes special data augmentations and performs extensive hyperparameter search to boost the performance of DeiT [62]. During the reproduction of DeiT3, we observe that it is sensitive to hyperparameters and prone to overfitting. Replacing the class token with GAP (global average pooling)[13] leads to a 0.7% top-1 accuracy drop for the DeiT3-Large model after 800 epochs of training. Therefore, we use the class token instead of GAP in the plain transformer and report the result in Table 2, where VisionLLaMA achieves a top-1 accuracy comparable to DeiT3. The detailed hyperparameter is listed in the appendix. Note that the accuracy on a single resolution does not provide comprehensive comparisons, we also evaluate the performance across different image resolutions as [13] and report the result in Table 11(supp.). As for DeiT3, we use the bicubic interpolation for the learnable positional encoding. Although these two models have comparable performance at the resolution of 224$\times$224, the gap is enlarged when the resolution is increased, which means our method generalizes better across different resolutions, which is a vital function for many downstream tasks such as object detection.
Pyramid Vision Transformer. We use the same architecture as Twins-SVT [12] and the detailed configuration is listed in Table 16 (supp.). We remove the conditional position encoding since VisionLLaMA already contains one kind of rotary position encoding. Therefore, VisionLLaMA is a convolution-free architecture. We do not tune the hyper-parameters and directly follow the setting provided in [12]. Although it’s suboptimal, it can still achieve competitive performance. As [12, 13], we do not use the class token and apply GAP. The result is shown in Table 2 and our method achieves comparable performance as Twins across various levels of models and outperforms Swin [42] consistently. We further compare the pyramid transformers using popular downstream tasks, which are shown in the later sections.
@@ 4.2.2 Self-Supervised Training
To make fair comparisons, we limit the training data to ImageNet-1K. We also exclude any component that utilizes CLIP [50], DALLE [51], or distillation, which can be orthogonally combined to further boost the performance. Our implementation is based on the MMPretrain framework [14]. We utilize the MAE framework and replace the encoder using VisionLLaMA while keeping other components unchanged. This minor modified setting forms a controlled experiment to evaluate the role of our approaches. Moreover, we use the same hyperparameter as [25], which is suboptimal to our method. Fortunately, this simple setting still achieves a significant performance boost over the strong baseline.
Full fine-tuning. In such a setting, the model is first initialized using the pre-trained weights and then trained for extra epochs with totally trainable parameters. Trained by 800 epochs on the ImageNet, VisionLLaMA-Base achieves 84.0% top-1 accuracy, which exceeds ViT-Base by 0.8%. Note that our method uses a mask ratio of 0.75 as [25], whose training speed is about 3 times faster than SimMIM [73]. We also increased the training epochs to 1600 to verify whether VisionLLaMA keeps the advantage given sufficient training resources. VisionLLaMA-Base achieves new state-of-art result among MAE variants, 84.3% top-1 accuracy, which outperforms ViT-Base by 0.9%. This result is even higher than MaskFeat [69] where new training objectives are proposed. Regarding full fine-tuning having a risk of performance saturation [41, 67], our boost is significant. Next, we resort to the linear probing metric to provide extra evaluations, which is considered a more reliable evaluation for representative learning by a recent work [9].
Table 3: Comparison with MIM SSL methods. $\dagger$: reproduced in MMPretrain.
Models Pretrain Epochs SFT Acc(%) LP Acc(%)
ViT-Base-MAE [25] 800 83.2 65.1
@@ 4.3.1 Supervised Training
Following [12, 42], we evaluate our method using semantic segmentation on the ADE20K [80] dataset. To make fair comparisons, we limit the baselines to only using ImageNet-1K in the pre-training stage. Specifically, we make use of the UperNet [72] framework and replace the backbone with pyramid VisionLLaMA. The detailed setting of the hyperparameter is shown in Section 0.D.7(supp). We report the result in Table 4(a). Our method outperforms both Swin and Twins by more than 1.2% mIoU.
Table 4: Segmentation results on ADE20k.
(a) Performance comparisons with different backbones on ADE20K validation dataset. All backbones are pre-trained on ImageNet-1K with labels.
Models Param mIoU (ss)
@@
(b) Performance comparisons with different SSL trained backbones on ADE20K validation dataset. All models are pre-trained on ImageNet-1K without labels.$\dagger$: reproduced using [15].
Models Epochs mIoU(ss)
ViT-B 800 46.2
@@
4.3.2 Self-Supervised Training
We use UperNet [72] to perform semantic segmentation on ADE20K. We carefully control the experiment and replace the ViT backbone with VisionLLaMA while keeping other components and hyperparameters unchanged. The detailed hyperparameters are provided in Section 0.D.6(supp.). The result is given in Table 4(b). As for the 800 epoch pre-training groups, VisionLLaMA-B significantly boosts ViT-Base by 2.8% mIoU. It also outperforms some other modifications such as introducing extra training objectives or features [41, 69] by clear margins. Moreover, those approaches introduce extra overhead for the training process and slow down the training speed. We emphasize that the training speed of a method is becoming more and more important in the age of large models. In contrast, VisionLLaMA only involves the replacement of the base model and has the same fast training speed as [25]. In principle, our method can be seamlessly combined with these modifications. We further evaluate the performance of longer pre-training epochs of 1600, VisionLLaMA-B achieves 50.2% mIoU on the ADE20K validation set, which boosts ViT-B by 2.1% mIoU.